Update from Erebus: First “real science” of the Season

Just a quick update before I go up to the crater to “supervise” the air-drop of all the drums of our new power cable. It’s blowing 20-25 knots at the moment, so we will have to wait and see if the pilot is willing to attempt it today or not.

The team lands at Fang camp ready for a few days of acclimatisation. The weather was perfect, and the views were stunning!

So, after a few days of being delayed at McMurdo due to bad weather, we made it into our acclimatisation camp at Fang Ridge last Friday. Saturday was thanksgiving, so we all walked up to LEH (Lower Erebus Hut – our main field camp) for dinner, then drove back to Fang in order to sleep at lower altitude again. Sunday morning, we all drove up to LEH and started unpacking. The conditions looked pretty good for spectroscopy and we managed to get a couple of instruments up and running. Unfortunately, there were too many ice crystals floating around in the air, and so we ended up recording negative gas amounts in the plume!

A vertically rising plume and no humidity made for a perfect spectroscopy day. Here and AvoScanner outside the garage at LEH scans horizontally accross the plume.

After that, we gave up on spectroscopy for a bit and concentrated on setting stuff up. The microwave ethernet link to the crater rim is up and running, as is the old crater rim power system. The thermal camera is almost ready to be deployed, just awaiting some final tests to make sure it doesn’t overheat in its box!

Aaron checks out the entrance to Kachina cave during our trip to Cones. This was the entrance that we used last year to access the cave, but this year it has completely closed up. Another entrance further downslope can still be used to get into the cave.

Yesterday was the first “real science” day. Conditions were perfect, with very little humidity and a vertically rising plume. After setting up an AvoScanner to measure Sulphur Dioxide amounts in the plume, I went out to Cones (the communications repeater station on the McMurdo side of Erebus) with Aaron to try and fix the seismic station data feed from LEH. We also met up with some Alt. Energy technicians from McMurdo to discuss power systems for LEH, Cones and the crater rim. After that, we teamed up with Clive and took the FTIR spectrometer into Warren Cave. The FTIR measures absorption of infra-red light by different gas species, and can identify the relative amounts of many different gas species. We are hoping to identify the species responsible for the hydrocarbon smell around the entrance to Warren – watch this space for the results!

Clive pours liquid Nitrogen into the FTIR to cool the detector during our trip to Warren cave. Carrying the (very fragile!) spectrometer and all the batteries, tripods, computers etc. into the cave was hard work - even though we were very close to the entrance.

In other news – the rest of the G081 team have all arrived safe and sound in McMurdo now and will be heading to Fang camp at the weekend. We expect to see them at LEH next Monday.

This entry was posted in Fieldwork, Mt. Erebus. Bookmark the permalink.

Leave a Reply

Your email address will not be published. Required fields are marked *

You may use these HTML tags and attributes: <a href="" title=""> <abbr title=""> <acronym title=""> <b> <blockquote cite=""> <cite> <code> <del datetime=""> <em> <i> <q cite=""> <strike> <strong>